OpenStudy (anonymous):

Solve \(\dfrac{a}{ax-1}\dfrac{b}{bx-1}\)=a+b by factorization. please solve it step by step by explaining the detail

OpenStudy (anonymous):

do u know the ans???

OpenStudy (cj49):

is der + sign between a/ax-1 nd b/bx-1

OpenStudy (anonymous):

@ganeshie8 @adilalvi @dan815 @iambatman @abhisar @mathsolver @phi

OpenStudy (anonymous):

@Sheraz12345 @paki @thomaster @chmvijay can you expain this

OpenStudy (bradely):

What symbol between a/ax-1and b/bx-1

OpenStudy (anonymous):

\(\dfrac{a}{ax-1}+\dfrac{b}{bx-1}\)=a+b i rpeat the question again sorry @cj49 i didn't understand what u mean

OpenStudy (anonymous):

@ash2326 @ParthKohli @sidsiddhartha can you help me

OpenStudy (anonymous):

@bradely what happen

OpenStudy (cj49):

a/(ax-1) + b/(bx-1) = -(a+b) Multiply by (ax-1)*(bx-1) a(bx -1) + b(ax-1) = -(a+b)(ax-1)*(bx-1) 2abx - (a + b) = -(a+b)(abx^2 - (a+b)x + 1) 2abx = -(a + b)abx^2 + x(a+b)^2 (a + b)abx^2 - (a^2 + b^2) x = 0 This is a quadratic and can be solved using the standard form for the solution: X = [-B +/- SQRT(B^2 - 4AC)] / 2A x = [ (a^2 + b^2) +/- SQRT((a^2 + b^2)^2) ] / [ 2(a + b)ab] x = [ (a^2 + b^2) +/- (a^2 + b^2 ] / [ 2(a + b)ab] x = 0 x = (a^2 + b^2) / [ (a + b)ab] Check: x= 0 yes. Check: x = (a^2 + b^2) / [ (a + b)ab] a/(ax - 1) = a/[(a^2 + b^2) / [ (a + b)b] - 1] a/(ax - 1) = a/[(a^2 + b^2 - (a + b)b) / [ (a + b)b] ] a/(ax - 1) = a/[(a^2 - ab) / [ (a + b)b] ] a/(ax - 1) = b(a + b)/[(a - b) ] b/(bx - 1) = b/[(a^2 + b^2) / [ (a + b)a] - 1] b/(bx - 1) = a(a + b) / [b - a] b(a + b)/[(a - b) ] + a(a + b) / [b - a] (a - b)(a + b)/(b - a) = -(a + b) -(a + b) = -(a + b)

OpenStudy (anonymous):

in 3rd line -(a+b)(abx^2-(a+b)x+1 i don't understand @cj49

OpenStudy (anonymous):

\((-a+b)(abx^2-(a+b)x+1\)) how inn mid a+b has come

OpenStudy (anonymous):

from your 3rd line or equation i can't understand @cj49

OpenStudy (cj49):

(ax-1)(bx-1)=abx2-ax-bx+1 taking out x as a common factor in -ax-bx abx2-(a+b)x+1

OpenStudy (anonymous):

thank you @cj49